# BroadGen: A Framework for Generating Effective and Efficient Advertiser Broad Match Keyphrase Recommendations

### v3

@@ IV-A2 Data Augmentation
The primary limitation of BroadGen’s methodology lies in its dependency on historical interaction data between items and queries for the generation of keyphrases. Although the algorithm functions effectively under warm start scenarios, it encounters difficulties with new or “cold” items that have limited interaction data available. To address this shortcoming, our approach incorporates input from state-of-the-art models deployed at eBay to provide precise keyphrase recommendations for cold-start items [3, 2], thereby bolstering the performance of our algorithm when dealing with cold-start conditions. [^6] This kind of data augmentation has been studied in [20].
@@
TABLE I: Details of meta categories from eBay.
@@ IV-A3 Datasets
We present results on 4 meta-categories of eBay, CAT_1 to CAT_4, each representing a group whose size is based on number of items and queries. The category sizes are largest from CAT_1 to CAT_4, the detailed statistics for each category is presented in Table I. For each category, each item’s future query distribution $Q$ is limited to 1000 queries. Each of these 1000 queries has at least 1 token overlap with the item’s title.
@@
TABLE II: Precision, recall and F1 metrics for all the models. Darker hues represent more favorable outcomes.
@@ IV-B1 Defining Relevance
In eBay’s Advertising ecosystem, a downstream relevance model filters the queries matched by broad match keyphrases, which is done by eBay Search. Hence, it is crucial to avoid unnecessary recommendations that are too generic which impacts the downstream efficiency and focus on passing the subsequent filtering processes by Search on the matched queries. In addition, the queries have to be relevant to the item in terms of human judgment, or else it will not be engaged by the buyers. To assess the relevance between item and query, we employ a BERT [44] cross-encoder model [45] that has been fine-tuned using judgment data derived from Large Language Models (LLMs) at eBay [12]. This model’s relevance judgment emulates human judgment while still aligning with eBay Search’s definition of relevance. We use this relevance judgment to check whether a query is relevant to an item for our evaluation. Employing this cross-encoder’s judgment also captures the semantic nuances of relevance that traditional lexical scoring methods generally overlook [8].
@@ IV-B2 Relevant Reach Estimation (RRE)
The relevant reach of a broad match keyphrase is subject to the distribution of buyers’ search queries, the broad matching mechanism and sellers’ advertising bids and budget. Although it’s infeasible to compute this metric offline, we can approximate it through sampling from the Post-query distribution. In addition, we define a relevance function $\hat{R}(i,q)$ based on the definition relevance explicated in IV-B1 and a broad match function B(k, q) as followed:
@@ IV-D1 Dependency on Input Queries
TABLE IV: Effect of number of pre-queries on BroadGen.
@@
Figure 5: Comparison of Precision, Recall and F1 for BroadGen(Aug) across keyword counts and categories.
@@ IV-D2 Various Clustering Algorithms
Different clustering algorithms can replace the BroadGen clustering core outlined in Section III-C. Algorithms like HDBSCAN [46] are unsuitable for BroadGen due to its requirement for a minimum cluster size, which can lead to grouping dissimilar queries. Table V compares BroadGen’s performance when its agglomerative core is replaced with Affinity Propagation [47] and Bayesian Gaussian Mixture [48]. We present F1 scores based on the RRE metric alongside PTR scores. The Agglomerative method shows superior F1 scores in CAT_1-3, albeit slightly lower in CAT_4. Its PTR scores are notably higher than the other algorithms, as it offers greater flexibility in clustering, enabling extraction of longer keyphrases beneficial for match-type keyword recommendations. KMeans and Spectral Clustering were also evaluated, but their excessive runtime of over 140140ms and 190190ms per input on average impedes daily batch processing.
@@ IV-E1 Production Implementation
The daily recommendations of BroadGen(Aug) is made possible through eBay’s PySpark-based [49] batch inference system. The speed of our model is considerable — with 1000 Spark executors and 4 cores, it only takes 45 minutes to finish the whole batch inference on a space of 1.5 billion items and their corresponding Queries. The inputs of the production model are pre-queries with at least 5 impressions in the past 30 days augmented with exact keyphrases that pass the Bert relevance filter [12] from two of our recall models — fastText [50, 51] and Graphite [3]. These two recall models are trained as Extreme Multi-Label Classifiers [52] with relevant past queries as labels for exact keyphrase recommendations. With BroadGen’s processing ability, we can incorporate newer queries on a daily basis accounting for shifting label distributions (see Section I-A1) thus solving Challenge I-B4.
@@ IV-E3 Sensitivity Analysis
Figure 6: Performance drift over time for the BroadGen keyphrases in the treatment group vs baseline exact match and broad match keyphrases from the control group
@@ V Conclusion and Future Work
We introduce a new framework for advancing non-exact match type recommendations, addressing the challenges associated with exact keyphrase suggestions in sponsored search advertising. Our distributed model, BroadGen, employs a string clustering approach to offer broad match recommendations, efficiently handling up to a billion items in under an hour. Without access to ground truth match type keyphrases, we construct keyphrases by modeling the underlying token correspondence in historical queries. We also propose a unique evaluation methodology and metrics to evaluate broad match performance when real-world impressions and clicks are not available. In the future, we aim to expand our approach by implementing neural network equivalents to the cores of our framework. Such as, using text embeddings for input matrix generation while using a k-means or similar clustering algorithm for grouping the queries. Text summarization for the cluster representation core can be tricky as they tend to generate too generic or specific keyphrases. With the adoption of more complex cores based on LLMs we can subsequently cater to broader definitions of semantic broad match [26, 27, 53] and extend beyond token-based broad match in the future. Training different aspects of such a model will be difficult due to the absence of concrete supervised signals and the constraints of broad/phrase match recommendation. Moreover, the latency and infrastructure costs of these models will also be a concern due to growing number of items and queries.

### v5

@@ IV-A2 Data Augmentation
The primary limitation of BroadGen’s methodology lies in its dependency on historical interaction data between items and queries for the generation of keyphrases. Although the algorithm functions effectively under warm start scenarios, it encounters difficulties with new or “cold” items that have limited interaction data available. To address this shortcoming, our approach incorporates input from state-of-the-art models deployed at eBay to provide precise keyphrase recommendations for cold-start items [ashirbad-etal-2024, mishra2025graphexgraphbasedextractionmethod], thereby bolstering the performance of our algorithm when dealing with cold-start conditions.
@@
TABLE I: Number of items, queries and average queries per item in product categories from eBay.
@@ IV-A3 Datasets
We present results on 4 meta-categories of eBay, CAT_1 to CAT_4, each representing a group whose size is based on number of items and queries. The category sizes are largest from CAT_1 to CAT_4, the detailed statistics for each category is presented in Table I. For each category, a buyer search query is associated with an item if the buyer clicked on it (see curated datasets in Graphite [ashirbad-etal-2024]). Each item’s future query distribution $Q$ is limited to 1000 aggregated queries. Each of these 1000 queries has at least 1 token overlap with the item’s title.
@@
TABLE II: Precision, recall and F1 scores (RRE) for all the models. Darker hues represent more favorable outcomes.
@@ IV-B1 Defining Relevance
In eBay’s Advertising ecosystem, a downstream relevance model filters the queries matched by broad match keyphrases, which is done by eBay Search. Hence, it is crucial to avoid unnecessary recommendations that are too generic which impacts the downstream efficiency and focus on passing the subsequent filtering processes by Search on the matched queries. In addition, the queries have to be relevant to the item in terms of human judgment, or else it will not be engaged by the buyers. To assess the relevance between item and query, we employ a BERT [devlin-etal-2019-bert] cross-encoder model [wolf-etal-2020-transformers] that has been fine-tuned using judgment data derived from Large Language Models (LLMs) at eBay [dey2025judgejudgeusingllm]. This model’s relevance judgment emulates human judgment while still aligning with eBay Search’s definition of relevance. We use this relevance judgment to check whether a query is relevant to an item for our evaluation. Employing this cross-encoder’s judgment also captures the semantic nuances of relevance that traditional lexical scoring methods generally overlook [dey2025middlemanbiasadvertisingaligning].
@@ IV-B2 Relevant Reach Estimation (RRE)
The relevant reach of a broad match keyphrase is subject to the distribution of buyers’ search queries, the broad matching mechanism, and sellers’ advertising bids and budget. Although it’s infeasible to compute this metric offline, we can approximate it through sampling from the Post-query distribution. In addition, we define a relevance function $\hat{R}(i,q)$ based on the definition of relevance explicated in IV-B1 and a broad match function B(k, q) as follows:
@@ IV-D1 Dependency on Input Queries
TABLE IV: Effect of number of input pre-queries on RRE metrics for BroadGen.
@@
Figure 5: Comparing Precision, Recall and F1 for BroadGen(Aug) with increasing count of output keywords across the categories.
@@ IV-D2 Various Clustering Algorithms
Different clustering algorithms can replace the BroadGen clustering core outlined in Section III-C. Algorithms like HDBSCAN [hdbscan] are unsuitable for BroadGen due to its requirement for a minimum cluster size, which can lead to grouping dissimilar queries. Table V compares BroadGen’s performance when its agglomerative core is replaced with Affinity Propagation [affinityPropagation] and Bayesian Gaussian Mixture [bishop-2007]. We present F1 scores based on the RRE metric alongside PTR scores. The Agglomerative method shows superior F1 scores in CAT_1-3, albeit slightly lower in CAT_4. Its PTR scores are notably higher than the other algorithms, as it offers greater flexibility in clustering, enabling extraction of longer keyphrases beneficial for match-type keyword recommendations. KMeans and Spectral Clustering were also evaluated, but their excessive runtime of over 140140ms and 190190ms per input on average impedes daily batch processing.
@@ IV-E1 Production Implementation
The daily recommendations of BroadGen(Aug) is made possible through eBay’s PySpark-based [apache_spark] batch inference system. The speed of our model is considerable — with 1000 Spark executors and 4 cores, it only takes 45 minutes to finish the whole batch inference on a space of 1.5 billion items and their corresponding Queries. The inputs of the production model are pre-queries with at least 5 impressions in the past 30 days augmented with exact keyphrases that pass the Bert relevance filter [dey2025judgejudgeusingllm] from two of the recall models — fastText [fastText1, fastText2] and Graphite [ashirbad-etal-2024]. These two recall models are trained as Extreme Multi-Label Classifiers [xmc] with relevant past queries as labels for exact keyphrase recommendations. With BroadGen’s processing ability, we can incorporate newer queries on a daily basis accounting for shifting label distributions (see Section I-A1) thus solving Challenge I-B4.
@@ IV-E3 Sensitivity Analysis
Figure 6: Performance drift over time for the BroadGen keyphrases in the treatment group vs baseline keyphrases (both exact and broad match types) from the control group.
@@ V Conclusion and Future Work
We introduce a new framework for advancing non-exact match type recommendations, addressing the challenges associated with exact keyphrase suggestions in sponsored search advertising. Our distributed model, BroadGen, employs a string clustering approach to offer broad match recommendations, efficiently handling up to a billion items in under an hour. Without access to ground truth match type keyphrases, we construct keyphrases by modeling the underlying token correspondence in historical queries. We also propose a unique evaluation methodology and metrics to evaluate broad match performance when real-world impressions and clicks are not available. In the future, we aim to expand our approach by implementing neural network equivalents to the cores of our framework. Such as, using text embeddings for input matrix generation while using a k-means or similar clustering algorithm for grouping the queries. Text summarization for the cluster representation core can be tricky as they tend to generate too generic or specific keyphrases. With the adoption of more complex cores based on LLMs we can subsequently cater to broader definitions of semantic broad match [ebayMatchTypes, googleMatchTypes, semantic_approach_sigir07] and extend beyond token-based broad match in the future. Training different aspects of such a model will be difficult due to the absence of concrete supervised signals and the constraints of broad/phrase match recommendation. Moreover, the latency and infrastructure costs of these models will also be a concern due to growing number of items and queries.